OpenStudy (anonymous):

If the length of side a is 16 centimeters, the length of side b is 10 centimeters, and m∠A = 42°, what is the measure of ∠B ? Round your answer to two decimal places.

OpenStudy (anonymous):

This is a triangle right? Do you have a diagram or anything?

OpenStudy (anonymous):

yes its really just a regular triangle

OpenStudy (anonymous):

Then use law of sines \[\frac{ \sin42 }{ 16 }=\frac{ sinB }{ 10 }\]

OpenStudy (anonymous):

26.25?

OpenStudy (anonymous):

Hmm, that's close, but not what I got.

OpenStudy (anonymous):

okay what did you get?

OpenStudy (anonymous):

sin(42)/16=0.0418206628974375 0.0418206628974375*10=0.418206628974375 asin(0.418206628974375)=24.72141615429

OpenStudy (anonymous):

oh I was doing the problem like a porportion

OpenStudy (anonymous):

could you help me with some more?

OpenStudy (anonymous):

No, you just solve for B

OpenStudy (anonymous):

so could you ?